OpenStudy (anonymous):

12 trees she needs are currently in one location, 40 m from the spot the first tree will be planted. the trees will be spaced 6 m apart. the cart she uses can only transport one at a time. so she must take the first tree to the spot and then return to take another one and so on. after brittany takes all 12 trees to the correct spot and returns to the orginal location of the trees, how far has she travelled in total? arithmetic sequence or series? or geometric? and how to solve?

OpenStudy (anonymous):

it's arithmetic series if that helps out..

OpenStudy (anonymous):

Taking 40m as the first term in the arithmetic sequence. The distance to n tree= 40+ (n-1) 6 Total distance she walked is 2 times the sum of all the terms from n=1 to n = 12. Total distance = 2* (12/2) ( 2 * 40 + (12-1) * 6) = 12 (80 + 66) =1752